OpenStudy (anonymous):

Last week, Eva spent seven and three-fifths hours on homework. This week, she spent ten and six-tenths hours. How many more hours did she spend this week? 2 two and three-fifths 3 three and three-fifths

OpenStudy (tester97):

\[10\frac{ 6 }{ 10 }-7\frac{ 3 }{ 5 }\] Convert to LCD \[10\frac{ 6 }{ 10 }-7\frac{ 6 }{ 10 }\] Now since they are both 6/10 they cancel out \[10-7=3\] So that is your answer~